Find the derrivative of
f(x) = tan⁻¹((1-x)/(1+x))
1
Expert's answer
2020-07-01T17:55:43-0400

X=tanθ

Y=tan-1((1-tanθ )/(1+tanθ))

=tan-1((tan(π /4)-tan θ )/(1+tan( "\pi"  /4)tanθ ))

=tan-1(tan("\pi"/4-θ))

="\pi" /4-θ,

Formula (tan(π /4))=1,

dy/dx=d/dx("\pi" /4-tan-1x)

Answer: dy/dx=-1/(1+x2)